OpenStudy (anonymous):

a1=5 an=3an-l+2 what is the third item?

OpenStudy (aum):

\[ \text{The sequence is: }a_1, ~~a_2,~~ a_3,~~ ..., ~~a_n \\ a_1 = 5 \\ a_n = a_{n-1} + 2 \\ \text{To find second term, put n = 2:} \\ a_2 = a_1 + 2 = 5 + 2 = 7 \\ \text{To find third term, put n = 3:} \\ a_3 = a_2 + 2 = 7 + 2 = 9 \\ \text{The third term is: }9. \]